OpenStudy (anonymous):

what is the integral of 2/ (2x-x^2)^1/2 my guesses is sin^-1 (whatever) +c maybe ?

OpenStudy (experimentx):

??

OpenStudy (experimentx):

integral of ??

OpenStudy (anonymous):

the integral of x/ (2x-x^2)^1/2

OpenStudy (experimentx):

try completing squares .. 2x - x^2 = 1 - (1-x)^2

OpenStudy (experimentx):

and let 1 - x = sin u, x = 1 + sin u

OpenStudy (experimentx):

not sure though ... haven't checked!!

OpenStudy (anonymous):

the complete square u gave me is wrong its 1-(x-1)^2 i guess right ?

OpenStudy (anonymous):

@quakeash , would you mind writing your question on drawing board again? Click the draw button below and rewrite your question on that

OpenStudy (anonymous):

 Is this your question ??

OpenStudy (anonymous):

exactly

OpenStudy (experimentx):

LOL ... it wasn't trig substitution, http://www.wolframalpha.com/input/?i=integrate+x%2F%28sqrt%282x+-+x^2%29%29

OpenStudy (anonymous):

It is pretty simple. First follow what @experimentX said